# GenTorrent: Scaling Large Language Model Serving with An Overley Network

### v2

@@ 3.1 Overview
The user nodes form a dynamic overlay network. Upon joining GenTorrent, a user node registers its IP address and public key with the verification nodes, whose IP addresses and public keys are public information. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [8]. For scalability in large deployments, verification nodes may divide the whole system into multiple regions and create a list of users and model nodes for each region. The requirement is that the number of users in the region should be sufficiently large to hide the requester’s identity, for example, $>1000$ users, and the number of model nodes should be sufficiently large to achieve load balancing. GenTorrent allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
The committee of verification nodes uses a recent BFT consensus protocol [7] to ensure the information maintained by them is consistent, assuming $<1/3$ of the nodes are malicious. The committee maintains a list of user nodes registered to GenTorrent, including their public keys, and a list of model nodes in GenTorrent, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes through the anonymous overlay and update their scores based on the responses.
@@ 3.2 Anonymous Overlay of Users
Using S-IDA. We briefly present an existing protocol Secure Information Dispersal Algorithm (S-IDA) [23] that is essential to achieve message confidentiality, anonymous routing, and failure resilience in GenTorrent. An Information Dispersal Algorithm (IDA) splits a message $M$ into multiple fragments, each of which is in ciphertext. If the receiver gets enough fragments, it can recover the plaintext message. Two most well-known IDAs are Shamir’s secret sharing (SSS) [38] and Rabin’s IDA [35]. Both of them requires $k$ out of $n$ fragments to recover the original message $M$. However, each has its weakness. SSS ensures any $k-1$ fragments provide no information about $M$, but each fragment has the same length as $M$, results in $n\times$ bandwidth cost. In Rabin’s IDA, each fragment is around $l/k$ long where $l$ is the size of $M$ but Rabin’s IDA does not guarantee that $<k$ fragments provide no information about $M$. Secure IDA (S-IDA) [23, 34] improves the security level of IDA by combining it with symmetric encryption, such as Advanced Encryption Standard (AES). The sender applies the following steps.
@@
Step 1. Preparation. When a user node $u$ starts running the GenTorrent client program, it joins the overlay network. The verification nodes’ IP and public keys are available to the public. The user $u$ contacts an arbitrary verification node to download a list of overlay users, called the user list, and a list of model nodes, called the mode node list. The user list is a consensus among all verification nodes and signed by more than 2/3 verification nodes. Each entry in the user list includes the public key and IP address of an overlay user. The users can be further grouped into regions, each of which has a separate list including the users around a certain network location. Users in the same region will be used to form anonymous paths. Grouping users into regions will significantly reduce the latency of anonymous paths, and a region should also include a sufficient number of users (e.g., over 1000) to ensure anonymity. In addition, GenTorrent must determine two parameters $n$ and $k$ ($n>k$) for $(n,k)$ S-IDA. It allows up to $n-k$ failed paths among $n$ overlay paths.
@@ 3.4 Model verification
$R(T)=\alpha\cdot R(T-1)+\beta\cdot C(T)\vspace{-1.5ex}$
where $R(T-1)$ is the score from the last epoch, $C(T)$ is the average score from the challenges in $T$, and $\alpha$ and $\beta$ are the weights for the moving average. In our implementation, we set $\alpha=0.4$ and $\beta=0.6$.
The punishment to the reputation for a low score should be much stronger than the reward for a high score. This ensures the model node has to respond correctly consistently to earn a high reputation. We use a sliding window approach to apply punishment to low values of $C(T)$. The verification nodes maintain a sliding window of each model including the past $W$ values of $C(T)$–the window size $W$ is set to 5 in our implementation. If $C(T)$ is smaller than a threshold $\tau$, it is considered an abnormal value. Let the count of abnormal values in the past $W$ be $c$. If $c/w$ is above a predefined threshold $\gamma$ (such as 1/5), the punishment applies as follows:
$R(T)=\alpha\cdot R(T-1)+\frac{W+1}{W+c/\gamma+2}\cdot C(T)\vspace{-1.5ex}$
@@
Verification is performed epoch-by-epoch. In each epoch, the committee collaboratively assesses the correctness of model nodes and updates their reputation accordingly. GenTorrent relies on a consensus protocol Tendermint [6]. Tendermint is an open-source protocol to provide BFT consensus that ensures safety and liveness under partial synchrony. The system assumes that out of $n$ nodes, a maximum of $f$ can be Byzantine faulty, where $n\leq 3f+1$. Each update message should be signed by at least $12n/3+$ 1 nodes before commitment. In each epoch, a verification node is selected as the leader to send challenge requests to model nodes. To reduce overwhelming model nodes, only the leader of each epoch communicates with the model nodes, collects their responses, and broadcasts the results to the committee. Each verification node then locally verifies the results independently and votes on the updates.
@@
If the response and digital signature do not match or the model node does not respond, the leader can include an “invalid response from model node $x$” message in the broadcast message, but does not reduce $x$’s reputation. It is to prevent the leader from attacking a model node by dropping or altering the response. Only when more than $1/3$ of the committee reported invalid responses for the same model node, implying that the model node misbehaves on purpose, $x$’s reputation will be reduced.
@@ 7 Conclusion
Strengths and weaknesses of slicing protocols. One idea is to use sliced messages for overlay anonymity [22, 34]. In garlic routing [34], the sender sends $n$ slices, each is about $1/k$ of the original message size long and is called a clove. If the receiver gets $k$ out of $n$ cloves, it can recover the original message. Garlic achieves 1) anonymity without encryption and decryption operations on the paths; 2) reliability under node failures; 3) the bandwidth cost does not increase significantly with the $n-k$ redundant cloves. Garlic routing has a few weaknesses if we apply it directly to GenTorrent. 1) Garlic relies on random walks to establish routing paths and the corresponding “proxies”. However, random walks add uncertainty to the success of finding enough proxies. 2) Garlic preserves the provider (model node) anonymity, which is unnecessary for GenTorrent.
@@
11:         if $node$ has child $chunk_{i}$ then
@@
5:     $\triangleright$ Cache hit
6:         $candidate\leftarrow$ Model node with fewest relative requests in $modelNodeList$
@@
4:     for $i\leftarrow 0$ to $length(output)-1$ do
@@
8:         if $token\_prob$ exists then
              $probabilities.append(token\_prob)$
@@
              $probabilities.append(\epsilon)$triangleright {Small constant}

### v3

# GenTorrent: Scaling Large Language Model Serving with An Overlay Network

@@ 3.1 Overview
The user nodes form a dynamic overlay network. Upon joining GenTorrent, a user node registers its IP address and public key with the verification nodes, whose IP addresses and public keys are public information. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [8]. For scalability in large deployments, verification nodes may divide the whole system into multiple regions and create a list of users and model nodes for each region. The requirement is that the number of users in the region should be sufficiently large to hide the requester’s identity, for example, >1000>1000 users, and the number of model nodes should be sufficiently large to achieve load balancing. GenTorrent allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
The committee of verification nodes uses a recent BFT consensus protocol [7] to ensure the information maintained by them is consistent, assuming <1/3<1/3 of the nodes are malicious. The committee maintains a list of user nodes registered to GenTorrent, including their public keys, and a list of model nodes in GenTorrent, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes through the anonymous overlay and update their scores based on the responses.
@@ 3.2 Anonymous Overlay of Users
Using S-IDA. We briefly present an existing protocol Secure Information Dispersal Algorithm (S-IDA) [23] that is essential to achieve message confidentiality, anonymous routing, and failure resilience in GenTorrent. An Information Dispersal Algorithm (IDA) splits a message $M$ into multiple fragments, each of which is in ciphertext. If the receiver gets enough fragments, it can recover the plaintext message. Two most well-known IDAs are Shamir’s secret sharing (SSS) [38] and Rabin’s IDA [35]. Both of them requires $k$ out of $n$ fragments to recover the original message $M$. However, each has its weakness. SSS ensures any k1k-1 fragments provide no information about $M$, but each fragment has the same length as $M$, results in $n\times$ bandwidth cost. In Rabin’s IDA, each fragment is around $l/k$ long where $l$ is the size of $M$ but Rabin’s IDA does not guarantee that <k<k fragments provide no information about $M$. Secure IDA (S-IDA) [23, 34] improves the security level of IDA by combining it with symmetric encryption, such as Advanced Encryption Standard (AES). The sender applies the following steps.
@@
Step 1. Preparation. When a user node $u$ starts running the GenTorrent client program, it joins the overlay network. The verification nodes’ IP and public keys are available to the public. The user $u$ contacts an arbitrary verification node to download a list of overlay users, called the user list, and a list of model nodes, called the mode node list. The user list is a consensus among all verification nodes and signed by more than 2/3 verification nodes. Each entry in the user list includes the public key and IP address of an overlay user. The users can be further grouped into regions, each of which has a separate list including the users around a certain network location. Users in the same region will be used to form anonymous paths. Grouping users into regions will significantly reduce the latency of anonymous paths, and a region should also include a sufficient number of users (e.g., over 1000) to ensure anonymity. In addition, GenTorrent must determine two parameters $n$ and $k$ ($n>k$) for $(n,k)$ S-IDA. It allows up to nkn-k failed paths among $n$ overlay paths.
@@ 3.4 Model verification
$R(T)=\alpha\cdot R(T-1)+\beta\cdot C(T)\vskip-6.45831pt$
where R(T1)R(T-1) is the score from the last epoch, $C(T)$ is the average score from the challenges in $T$, and $\alpha$ and $\beta$ are the weights for the moving average. In our implementation, we set $\alpha=0.4$ and $\beta=0.6$.
The punishment to the reputation for a low score should be much stronger than the reward for a high score. This ensures the model node has to respond correctly consistently to earn a high reputation. We use a sliding window approach to apply punishment to low values of $C(T)$. The verification nodes maintain a sliding window of each model including the past $W$ values of C(T)C(T)–the window size $W$ is set to 5 in our implementation. If $C(T)$ is smaller than a threshold $\tau$, it is considered an abnormal value. Let the count of abnormal values in the past $W$ be $c$. If $c/w$ is above a predefined threshold $\gamma$ (such as 1/5), the punishment applies as follows:
$R(T)=\alpha\cdot R(T-1)+\frac{W+1}{W+c/\gamma+2}\cdot C(T)\vskip-6.45831pt$
@@
Verification is performed epoch-by-epoch. In each epoch, the committee collaboratively assesses the correctness of model nodes and updates their reputation accordingly. GenTorrent relies on a consensus protocol Tendermint [6]. Tendermint is an open-source protocol to provide BFT consensus that ensures safety and liveness under partial synchrony. The system assumes that out of $n$ nodes, a maximum of $f$ can be Byzantine faulty, where $n\leq 3f+1$. Each update message should be signed by at least $2n/3+1$ nodes before commitment. In each epoch, a verification node is selected as the leader to send challenge requests to model nodes. To reduce overwhelming model nodes, only the leader of each epoch communicates with the model nodes, collects their responses, and broadcasts the results to the committee. Each verification node then locally verifies the results independently and votes on the updates.
@@
If the response and digital signature do not match or the model node does not respond, the leader can include an “invalid response from model node xx” message in the broadcast message, but does not reduce $x$’s reputation. It is to prevent the leader from attacking a model node by dropping or altering the response. Only when more than $1/3$ of the committee reported invalid responses for the same model node, implying that the model node misbehaves on purpose, $x$’s reputation will be reduced.
@@ 7 Conclusion
Strengths and weaknesses of slicing protocols. One idea is to use sliced messages for overlay anonymity [22, 34]. In garlic routing [34], the sender sends $n$ slices, each is about $1/k$ of the original message size long and is called a clove. If the receiver gets $k$ out of $n$ cloves, it can recover the original message. Garlic achieves 1) anonymity without encryption and decryption operations on the paths; 2) reliability under node failures; 3) the bandwidth cost does not increase significantly with the nkn-k redundant cloves. Garlic routing has a few weaknesses if we apply it directly to GenTorrent. 1) Garlic relies on random walks to establish routing paths and the corresponding “proxies”. However, random walks add uncertainty to the success of finding enough proxies. 2) Garlic preserves the provider (model node) anonymity, which is unnecessary for GenTorrent.
@@
11:   if nodenode has child $chunk_{i}$ then
@@
[…]
6:   $candidate\leftarrow$ Model node with fewest relative requests in modelNodeListmodelNodeList
@@
4:  for $i\leftarrow 0$ to length(output)1length(output)-1 do
@@
8:   if token_probtoken\_prob exists then
     probabilities.append(token_prob)probabilities.append(token\_prob)
@@
     $probabilities.append(\epsilon)$ $\triangleright$ {Small constant}